40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3692     Entry time: Mon Oct 11 22:04:28 2010
Author: yuta 
Type: Update 
Category: CDS 
Subject: damping for MCs are basically working 

Background:
 Even if we can't use the Dataviewer to get the time series of each 4 DOF displacements, we can still use StripTool to monitor the ringdowns and see if the damping servo is working.

What I did:
1. Excited vibration by kicking the mirror randomly (by putting some offsets randomly and turing the filters on and off randomly).

2. Turned all the servo off by clicking "ShutDown" button.

3. Turned all the servo on by clicking "Normal" button.

3. Monitored each 4 DOF displacements with StripTool and see if there are any considerably low-Q ringdown after turning on the servo.
 The values I monitored are as follows;
  C1:SUS_MCX_SUSPOS_INMON
  C1:SUS_MCX_SUSPIT_INMON
  C1:SUS_MCX_SUSYAW_INMON
  C1:SUS_MCX_SDSEN_INMON  (X=1,2,3)

All the settings I used for this observation are automatically saved here;
  /cvs/cds/caltech/burt/autoburt/snapshots/2010/Oct/11/21:07/c1mcs.epics

Result:
 Attached is the screenshots of StripTool Graph window for each 3 MCs.
 As you can see, the dampings for each DOF, each MCs are basically working.

Note:
 Do NOT turn off all the damping servo by clicking "Damp" buttons or setting the SUSXXX_GAIN to 0. It crashes c1mcs.

Next work:
 - check and relate the signal sign with the actual moving direction of the optics
 - fix data aquisition system
 - measure Q-values when the servo is on and off using DAQ and Dataviewer

Attachment 1: SUS-MC1.png  29 kB  | Hide | Hide all
SUS-MC1.png
Attachment 2: SUS-MC2.png  28 kB  Uploaded Mon Oct 11 23:04:49 2010  | Hide | Hide all
SUS-MC2.png
Attachment 3: SUS-MC3.png  30 kB  Uploaded Mon Oct 11 23:05:06 2010  | Hide | Hide all
SUS-MC3.png
ELOG V3.1.3-